Solve using basic differentiation rule


𝑔(π‘₯) =3βˆ’2π‘₯/3+2π‘₯

1
Expert's answer
2022-05-16T10:28:03-0400

"\frac{d}{dx}(3-\frac{2x}{3}+2x)=g’(x)"

"g’(x)=-\frac{2}{3}+2=\frac{4}{3}"